OpenStudy (anonymous):

\[u^3\sqrt(t)+ 2 \sqrt(t^3)\]

ganeshie8 (ganeshie8):

yes what are you trying to do ? :)

OpenStudy (anonymous):

derivative sorry

OpenStudy (anonymous):

Differentiate the sum term by term, this requires power rule and chain rule, \[u^3\frac{ d }{ dt }(\sqrt{t})+2\frac{ d }{ dt }\sqrt{t^3}\] I just factored the constants out for now, can you do this?

OpenStudy (anonymous):

It might be easier for you if you convert the square root to exponents \[\sqrt{t} = t ^{1/2}\]

OpenStudy (anonymous):

But hey, I'm going to sleep now, @ganeshie8 mind taking over :3?

OpenStudy (anonymous):

wait u^3 isnt a constant isnt it? u^3 1/2t^(-1/2) + 2 * 1/2t^(-1/2) * 3t^2

ganeshie8 (ganeshie8):

Is it given that \( u\) is a function of \(t\) ?

ganeshie8 (ganeshie8):

you're right if \(u\) is a constant, otherwise simply use chain rule for the first product

ganeshie8 (ganeshie8):

product rule* \[\large \left(fg\right)' = f'g + fg'\]

OpenStudy (anonymous):

oh, im sorry. i did the question wrong i still need help but instead of u to the 3rd its a different problem let me rewrite it

OpenStudy (anonymous):

\[u=\sqrt[3]{t^2}+2\sqrt{t^3}\]

ganeshie8 (ganeshie8):

let me rephrase the problem : \(\large f(u,t) = u^3\sqrt{t}+ 2 \sqrt{t^3}\) \(\large u=\sqrt[3]{t^2}+2\sqrt{t^3}\) \(\large \dfrac{df}{dt} = ?\)

ganeshie8 (ganeshie8):

is that the problem ? or have i mixed both the problems :o

OpenStudy (anonymous):

no you mixed both problems, the only problem is the u= cuberoot problem the last one i posted im sorry if i confused you

ganeshie8 (ganeshie8):

haha i thought so ! so you just want to find \(\large \dfrac{du}{dt}\) given this : \(\large u=\sqrt[3]{t^2}+2\sqrt{t^3}\) ?

OpenStudy (anonymous):

yes!

ganeshie8 (ganeshie8):

good, we're on same page finally :) so take the derivative, whats stopping you ?

ganeshie8 (ganeshie8):

\(\large u=\sqrt[3]{t^2}+2\sqrt{t^3}\) \(\large \dfrac{d}{dt}(u) = \dfrac{d}{dt} \left(\sqrt[3]{t^2}+2\sqrt{t^3}\right)\)

OpenStudy (anonymous):

i dont really understand how to do it. it becomes a mess. how would i go about taking the derivative of (t^2)^1/3 would you use the chain rule?

ganeshie8 (ganeshie8):

write \(\large \sqrt[3]{t^2}\) as \(\large t^{\frac{2}{3}}\) and use the power rule : \(\large \dfrac{d}{dx}(x^n) = nx^{n-1}\)

ganeshie8 (ganeshie8):

\(\large u=\sqrt[3]{t^2}+2\sqrt{t^3}\) \(\large \dfrac{d}{dt}(u) = \dfrac{d}{dt} \left(\sqrt[3]{t^2}+2\sqrt{t^3}\right)\) \(\large ~~~~~~~~~~~ = \dfrac{d}{dt} \left(t^{\frac{2}{3}}+2t^{\frac{3}{2}}\right)\)

OpenStudy (anonymous):

2/3t^(-1/3) + 3/2t^(1/3) ?

ganeshie8 (ganeshie8):

looks perfect !

ganeshie8 (ganeshie8):

wait a sec, who ate the \(2\) that was attached to the second term before differentiating ?

OpenStudy (anonymous):

i forgot the 2 lol it'd be the 2/3^(-1/3) + 2(3/2(^1/3)

ganeshie8 (ganeshie8):

yes ! we're done with the derivative part. calculus part is over... you may leave the answer as it is or simplify it - up to you :)

OpenStudy (anonymous):

i need help simplfying it thank you! i dont really understand how to simplify negative fractions or fractional exponents

ganeshie8 (ganeshie8):

sure, so where are we

ganeshie8 (ganeshie8):

\(\large ~~~~~~~~~~~ = \dfrac{d}{dt} \left(t^{\frac{2}{3}}+2t^{\frac{3}{2}}\right)\) \(\large ~~~~~~~~~~~ = \frac{2}{3}t^{-\frac{1}{3}} + 2\frac{3}{2}t^{\frac{1}{3}}\) \(\large ~~~~~~~~~~~ = \frac{2}{3}t^{-\frac{1}{3}} + 3t^{\frac{1}{3}}\)

ganeshie8 (ganeshie8):

fine so far ? :)

OpenStudy (anonymous):

mhm everythings perfect so far

ganeshie8 (ganeshie8):

somehow teachers don't like negative exponents, so we use this exponent rule : \[\large a^{-m} = \dfrac{1}{a^m}\]

ganeshie8 (ganeshie8):

\[\large ~~~~~~~~~~~ = \frac{2}{3}t^{-\frac{1}{3}} + 3t^{\frac{1}{3}}\] \[\large ~~~~~~~~~~~ = \frac{2}{3}\dfrac{1}{t^{\frac{1}{3}}} + 3t^{\frac{1}{3}}\]

ganeshie8 (ganeshie8):

\[\large ~~~~~~~~~~~ = \frac{2}{3\sqrt[3]{t}} + 3\sqrt[3]{t}\]

ganeshie8 (ganeshie8):

guess i would stop here..

OpenStudy (anonymous):

ohh is that really it? thank you very much youve been a huge help!

ganeshie8 (ganeshie8):

thats rly it ! yw :)